OpenStudy (anonymous):

21x^2-6=15x solve the equation check your solution show work

OpenStudy (anonymous):

21x^2-15x-6=0 factor 3 (x-1) (7 x+2) x = -2/7 x = 1

OpenStudy (anonymous):

If you want to solve problems faster during the SAT test math sections, try our SAT app for your calculator: http://download.cnet.com/SAT-Operating-System-For-TI-83-And-TI-84/3000-20415_4-75448635.html?part=dl-&subj=dl&tag=button

OpenStudy (anonymous):

i dont need to worry bout SAT testing yet im only a freshmen

OpenStudy (anonymous):

You can use it for the PSAT.

OpenStudy (anonymous):

and in your math class.

OpenStudy (amistre64):

i think you have stock in that site ;)

OpenStudy (anonymous):

You are correct. But we do try to help students on a one on one basis with math in general and through the use of our software.